# Multi-Objective Coverage Bayesian Optimization (MOCOBO)

### v2

@@ A.2.3 Wall-clock runtime of Algorithm 1
Table A.3: Average wall-clock runtime of Algorithm 1 for different values of $N$ (the number of data points), $K$ (the covering set size), and $T$ (the number of objectives). The “Runtime" column gives the average runtime of Algorithm 1 with parallel computation of the of the marginal coverage of adding the $N$ data points (parallel computation of lines 5-6). The “Runtime No Parallelization" column gives the average runtime of Algorithm 1 without this parallelization of the inner for loop. Average run times are computed by running Algorithm 1 on a NVIDIA RTX A5000 GPU and averaging over $10$ runs. Standard errors over the $10$ runs are also provided.
Efficiency of Algorithm 1
Algorithm 1 could also be made more efficient by pruning the points in $D_{s}$ with zero marginal coverage improvement after each outer-loop, since these points will continue to have zero marginal coverage improvement on subsequent loops. In future work, we plan to explore this and any other tricks that might allow us to further improve the efficiency of Algorithm 1.
@@ A.3 Approximation strategy for batch acquisition with Expected Coverage Improvement (ECI)
In batch acquisition, we select a batch of $q>1$ candidates for evaluation. In Equation 6, we define q-ECI, a natural extension of the ECI acquisition function defined in Equation 4 to the batch acquisition setting. q-ECI gives the expected improvement in the coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. When using batch acquisition, we aim to select a batch of $q$ points that maximize q-ECI.
@@
[…]
Instead of sampling $m$ batches of candidates, we sample $m$ individual data points $P=\left\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{m}\right\}$. For each sampled point $\mathbf{p}_{j}$, we sample a realization $\hat{\mathbf{y}}_{j}=(\hat{f}_{1}(\mathbf{p}_{j}),\ldots,\hat{f}_{T}(\mathbf{p}_{j}))$. As in Section 3.1.1, we use the sampled realizations $\hat{\mathbf{y}}_{j}$ to compute an approximate coverage improvement $CI(\mathbf{p}_{j})$ as defined in Equation 5 for each point $\mathbf{p}_{j}$. To obtain a batch of $q$ candidates, we then greedily select the $q$ points $\mathbf{p}_{j}\in P$ with the $q$ largest expected coverage improvements. Note that this strategy does not involve sequential optimization of the $q$ points, as the batch of $q$ points is selected simultaneously as the points with the top-$q$ expected coverage improvements.
@@ A.4 Limitations and future work
A primary limitation of MOCOBO is that the choice of $K$ (the covering set size) may not always be straightforward. For example, as we mention in Section 4, it may require domain knowledge to choose $K<T$ that is large enough that achieving good coverage is possible despite multiple conflicting objectives. In many practical applications, we prefer the smallest possible set size $K$ such that good coverage can still be achieved. In future work, we plan to explore methods for simultaneously optimizing both the solutions in the covering set, and the size of the covering set, balancing the trade-off between minimizing the number of solutions needed and maximizing overall coverage.
@@ A.6 Compute resources
We use GPU works to run all experiments and produce all empirical results provided in this paper. A single GPU was used per run of each method compared on each task. Each each run uses approximately 12-18 GB of the GPU memory. Most experiments were executed on our internal cluster of NVIDIA RTX A5000 GPUs (see internal cluster compute details in Table A.4). We also used cloud compute resources for two weeks to complete additional replications of some experiments. We used a total of eight RTX 4090 GPU workers from runpod.io, each with approximately 24 GB of GPU memory.
@@
Preliminary experiments refer to the initial experiments for e.g. method development that are not included as results in the paper. All preliminary experiments were done our internal cluster of NVIDIA RTX A5000 GPUs (see internal cluster compute details in Table A.4). We spent approximately $2000$ hours of GPU time on preliminary experiments.
@@ A.7 Additional implementation details
In this section, we provide additional implementation details for MOCOBO. After the paper is accepted, we will also refer readers to the MOCOBO codebase for the full-extent of implementation details and experimental setup needed to reproduce the results provided https://github.com/placeholder.
@@ A.7.2 Surrogate model
We train the PPGPR surrogate model(s) on data collected during optimization using the Adam optimizer with a learning rate of $0.001$ and a mini-batch size of $256$. On each step of optimization, we update the model on collected data until we stop making progress (loss stops decreasing for $3$ consecutive epochs), or exceed $30$ epochs. Since we collect a large amount of data for each optimization run (i.e. as many as $62e$ 6 data points in a single for the “template constrained" peptide design task), we avoid updating the model on all data collected at once. On each step of optimization, we update the current surrogate model only on a subset of $1000$ of the collected data points. This subset is constructed of the data that has obtained the highest objective values so far, along with the most recent batch of data collected. By always updating on the most recent batch of data collected, we ensure that the surrogate model is conditioned on every data point collected at some point during the optimization run.
@@ A.7.3 Initialization data
In this section, we provide details regarding the data used to initialize all optimization runs for all tasks in Section 4.
@@
To initialize optimization for the peptide design tasks, we generate a a set of $20,000$ peptide sequences by making random edits (insertions, deletions, and mutations) to the $10$ template peptide sequences in Table A.8. We generate labels for these $20,000$ peptides once, and then use the labeled data to initialize optimization for all methods compared.
@@ A.8 Additional task details
In this section we provide additional details for the chosen set of tasks we provide results for in Section 4.
@@ A.8.1 HDR image tone mapping
In Table A.5, we list the names of the $7$ image quality metrics used for the image tone mapping tasks described in Section 4.1.
@@
The target image quality (IQA) and aesthetic (IAA) metrics are organized in Table A.5, where all except nima are IAA metrics, while nima is an IQA metric. For more detailed information about each metric and their corresponding datasets, please refer to their original references.
@@
(The constants were taken from the code of Farbman et al. 18.) The luminance image $L$ is then logarithmically compressed and then decomposed into three detail layers and one base layer by applying the guided filter three times, each with a different radius parameter $r_{i}$ and smoothing parameter $\epsilon_{i}$ for $i=1,\ldots,3$. Then, we amplify or attenuate each channel with a corresponding gain coefficient $g_{\mathrm{detail},i}$ $i=1,\ldots,3$ for the detail and $g_{\mathrm{base}}$ for the base layers. The image is then reconstructed by adding all the layers, including the colors. Following Tumblin and Turk [62], we also apply a gain, $g_{\mathrm{color}}$, to the color channels. Finally, the resulting image is applied an overall gain $g_{\mathrm{out}}$ and then gamma-corrected [51, Section 2.9] with an exponent of $1/\gamma$. The parameters for this pipeline are organized in Table A.6. The implementation uses OpenCV [5], in particular, the guided filter implementation in the extended image processing (ximgproc) submodule.
@@
to the unit hypercube ${[0,1]}^{13}$. In particular, the mapped values on the unit interval $[0,1]$ linearly interpolate the domain of each parameter shown in Table A.6. For the radius parameters $r_{1},r_{2},r_{3}$, which are categorical, we naively quantize the domain by rounding the output of the interpolation to the nearest integer.
@@ A.8.2 Peptide design
Table A.7 specifies the $T=11$ target bacteria used for the peptide design task from Section 4. The first seven bacteria are Gram negative bacteria (Objective IDs B1-B7) and the last four (Objective IDs B8-B11) are Gram positive. Table A.8 gives the $10$ template amino acid sequences used for the “template constrained" variation of the peptide design task from Section 4.
@@ A.9 Proof that finding $S^{*}_{D_{s}}$ is NP-hard
In this section, we prove Lemma 3.1, that finding $S^{*}_{D_{s}}$ Is NP-Hard.
@@
We prove Lemma 3.1 this by reduction from the well-known Maximum Coverage Problem (MCP), which is NP-Hard.
@@
We do not claim the above proof of Lemma 3.1 as a novel contribution of this work, as it follows straightforwardly from the fact that the well-known Maximum Coverage Problem is NP-hard. Note that proving Lemma 3.1 would also follow straightforwardly from the work of Ding et al. [13] who demonstrate that the coverage optimization problem generalizes k-means clustering. Rather than providing additional novel contribution, the above proof of Lemma 3.1 serves to justify our use of a greedy approximation algorithm (Algorithm 1) to approximate $S^{*}_{D_{s}}$ on each iteration of MOCOBO.
A.10 Approximation proof for Algorithm 1
In this section, we prove Theorem 3.2, that Algorithm 1 is a $(1-\frac{1}{e})$-Approximation.
@@ A.10 Approximation proof for Algorithm 1
It follows that from Definition A.5 that:
@@
The coverage score $c(S)$ as defined in Definition A.3 is monotone, i.e., for any $S\subseteq S^{\prime}\subseteq D_{s}$, we have:
@@
The coverage score $c(S)$ as defined in Definition A.3 is submodular, i.e., for any $S\subseteq S^{\prime}\subseteq D_{s}$ and any $\mathbf{x}^{*}\in D_{s}\setminus S^{\prime}$, we have:
@@
Proof of Theorem 3.2 From Lemma A.6 and Lemma A.6, we know that $c(S)$ is monotone submodular.
Since $c(S)$ is monotone submodular, and Algorithm 1 approximates $A^{*}_{D_{s}}$ using greedy submodular optimization, it follows from Lemma A.8 that Algorithm 1 achieves:
@@ 3. Answer: [Yes]
  4. Justification: All stated claims are backed-up with results in Section 4 and the stated focus/scope of the paper accurately reflects what is discussed throughout the rest of the paper.
@@ 8. Answer: [Yes]
  9. Justification: See Section A.4 for discussion of limitations.
@@ 13. Answer: [Yes]
  14. Justification: All theoretical claims can be found in Section 3.1.2. All sets of assumptions and full proofs of each claim are provided in Section 3.1.2, Section A.9, and Section A.10. As far as we are aware, each proof is complete and correct.
@@ 18. Answer: [Yes]
  19. Justification: We provide detailed explanation of how our method works in Section 3 and all additional required details to reproduce results in Section 4 and Section A.7. Additionally, once the paper is accepted, we will replace the placeholder GitHub links with a link to a public GitHub repository containing the source code used to produce the results provided in Section 4. This source code will allow any reader to run our code to reproduce results on all tasks in the paper. Additionally, the README in the repository provides detailed instructions to make setting up the proper environment and running the code easy for users.
@@ 23. Answer: [Yes]
  24. Justification: Once the paper is accepted, we will replace the placeholder GitHub links in the main text with a link to a public GitHub repository containing the source code used to produce the results provided in this paper. This GitHub repository contains organized code that will allow any reader to run MOCOBO on all tasks to reproduce the results provided in this paper. Additionally, the README in the repository provides detailed instructions to make setting up the proper environment and running the code easy for users.
@@ 28. Answer: [Yes]
  29. Justification: All chosen hyper-parameters and implementation details are stated in section 4 and Section A.7.
@@ 38. Answer: [Yes]
  39. Justification: All compute details are provided in Section A.6.
@@ 48. Answer: [Yes]
  49. Justification: See discussion of broader societal impacts in Section A.5.
@@ 58. Answer: [Yes]
  59. Justification: Creators of all assets used to produce all results in this paper are cited in Section 4. All assets used are open source software or models.

### v3

@@ A.2.3 Wall-clock runtime of Algorithm˜1
Table A.3: Average wall-clock runtime of Algorithm˜1 for different values of $N$ (the number of data points), $K$ (the covering set size), and $T$ (the number of objectives). The “Runtime" column gives the average runtime of Algorithm˜1 with parallel computation of the of the marginal coverage of adding the $N$ data points (parallel computation of lines 5-6). The “Runtime No Parallelization" column gives the average runtime of Algorithm˜1 without this parallelization of the inner for loop. Average run times are computed by running Algorithm˜1 on a NVIDIA RTX A5000 GPU and averaging over $10$ runs. Standard errors over the $10$ runs are also provided.
Efficiency of Algorithm˜1
Algorithm˜1 could also be made more efficient by pruning the points in $D_{s}$ with zero marginal coverage improvement after each outer-loop, since these points will continue to have zero marginal coverage improvement on subsequent loops. In future work, we plan to explore this and any other tricks that might allow us to further improve the efficiency of Algorithm˜1.
@@ A.3 Approximation strategy for batch acquisition with Expected Coverage Improvement (ECI)
In batch acquisition, we select a batch of $q>1$ candidates for evaluation. In Equation˜6, we define q-ECI, a natural extension of the ECI acquisition function defined in Equation˜4 to the batch acquisition setting. q-ECI gives the expected improvement in the coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. When using batch acquisition, we aim to select a batch of $q$ points that maximize q-ECI.
@@
Here, $c(S^{*}_{D_{s}\cup\left\{(\mathbf{b}_{jr},\hat{\mathbf{y}}_{jr})\right\}})$ is the approximation of the coverage score of the new best covering set if we choose to evaluate candidate $\mathbf{b}_{jr}$, assuming the candidate point will have the sampled objective values $\hat{\mathbf{y}}_{jr}$. We would like to select and evaluate the batch of candidates $\mathbf{B}_{j}$ with the largest $q\text{-CI}$.
Evaluating $q\text{-CI}$ for a single candidate batch requires $q$ evaluations of $c(S^{*}_{D_{s}\cup\left\{(\mathbf{b}_{jr},\hat{\mathbf{y}}_{jr})\right\}})$. Each evaluation of $c(S^{*}_{D_{s}\cup\left\{(\mathbf{b}_{jr},\hat{\mathbf{y}}_{jr})\right\}})$ requires a call to Algorithm˜1 to first construct $S^{*}_{D_{s}\cup\left\{(\mathbf{b}_{jr},\hat{\mathbf{y}}_{jr})\right\}}$. Thus, batch acquisition with a full MC approximation of q-ECI requires $O(q\times m)$ calls of Algorithm˜1. Assuming a sufficiently large $m$ to achieve a reliable MC approximation, this can become expensive for large batch sizes $q$. We therefore propose a faster approximation of batch ECI for practical use with large $q$.
Instead of sampling $m$ batches of candidates, we sample $m$ individual data points $P=\left\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{m}\right\}$. For each sampled point $\mathbf{p}_{j}$, we sample a realization $\hat{\mathbf{y}}_{j}=(\hat{f}_{1}(\mathbf{p}_{j}),\ldots,\hat{f}_{T}(\mathbf{p}_{j}))$. As in Section˜3.1.1, we use the sampled realizations $\hat{\mathbf{y}}_{j}$ to compute an approximate coverage improvement $CI(\mathbf{p}_{j})$ as defined in Equation˜5 for each point $\mathbf{p}_{j}$. To obtain a batch of $q$ candidates, we then greedily select the $q$ points $\mathbf{p}_{j}\in P$ with the $q$ largest expected coverage improvements. Note that this strategy does not involve sequential optimization of the $q$ points, as the batch of $q$ points is selected simultaneously as the points with the top-$q$ expected coverage improvements.
@@ A.4 Limitations and future work
A primary limitation of MOCOBO is that the choice of $K$ (the covering set size) may not always be straightforward. For example, as we mention in Section˜4, it may require domain knowledge to choose $K<T$ that is large enough that achieving good coverage is possible despite multiple conflicting objectives. In many practical applications, we prefer the smallest possible set size $K$ such that good coverage can still be achieved. In future work, we plan to explore methods for simultaneously optimizing both the solutions in the covering set, and the size of the covering set, balancing the trade-off between minimizing the number of solutions needed and maximizing overall coverage.
@@ A.6 Compute resources
We use GPU works to run all experiments and produce all empirical results provided in this paper. A single GPU was used per run of each method compared on each task. Each each run uses approximately 12-18 GB of the GPU memory. Most experiments were executed on our internal cluster of NVIDIA RTX A5000 GPUs (see internal cluster compute details in Table˜A.4). We also used cloud compute resources for two weeks to complete additional replications of some experiments. We used a total of eight RTX 4090 GPU workers from runpod.io, each with approximately 24 GB of GPU memory.
@@
Preliminary experiments refer to the initial experiments for e.g. method development that are not included as results in the paper. All preliminary experiments were done our internal cluster of NVIDIA RTX A5000 GPUs (see internal cluster compute details in Table˜A.4). We spent approximately $2000$ hours of GPU time on preliminary experiments.
@@ A.7 Additional implementation details
In this section, we provide additional implementation details for MOCOBO. We also refer readers to the MOCOBO codebase for the full-extent of implementation details and experimental setup needed to reproduce results provided https://github.com/nataliemaus/mocobo.
@@ A.7.2 Surrogate model
We train the PPGPR surrogate model(s) on data collected during optimization using the Adam optimizer with a learning rate of $0.001$ and a mini-batch size of $256$. On each step of optimization, we update the model on collected data until we stop making progress (loss stops decreasing for $3$ consecutive epochs), or exceed $30$ epochs. Since we collect a large amount of data for each optimization run (i.e. as many as $2e6$ data points in a single for the “template constrained" peptide design task), we avoid updating the model on all data collected at once. On each step of optimization, we update the current surrogate model only on a subset of $1000$ of the collected data points. This subset is constructed of the data that has obtained the highest objective values so far, along with the most recent batch of data collected. By always updating on the most recent batch of data collected, we ensure that the surrogate model is conditioned on every data point collected at some point during the optimization run.
@@ A.7.3 Initialization data
In this section, we provide details regarding the data used to initialize all optimization runs for all tasks in Section˜4.
@@
To initialize optimization for the peptide design tasks, we generate a a set of $20,000$ peptide sequences by making random edits (insertions, deletions, and mutations) to the $10$ template peptide sequences in Table˜A.8. We generate labels for these $20,000$ peptides once, and then use the labeled data to initialize optimization for all methods compared.
@@ A.8 Additional task details
In this section we provide additional details for the chosen set of tasks we provide results for in Section˜4.
@@ A.8.1 HDR image tone mapping
In Table˜A.5, we list the names of the $7$ image quality metrics used for the image tone mapping tasks described in Section˜4.1.
@@
The target image quality (IQA) and aesthetic (IAA) metrics are organized in Table˜A.5, where all except nima are IAA metrics, while nima is an IQA metric. For more detailed information about each metric and their corresponding datasets, please refer to their original references.
@@
(The constants were taken from the code of Farbman et al. 18.) The luminance image $L$ is then logarithmically compressed and then decomposed into three detail layers and one base layer by applying the guided filter three times, each with a different radius parameter $r_{i}$ and smoothing parameter $\epsilon_{i}$ for $i=1,\ldots,3$. Then, we amplify or attenuate each channel with a corresponding gain coefficient $g_{\mathrm{detail},i}$ $i=1,\ldots,3$ for the detail and $g_{\mathrm{base}}$ for the base layers. The image is then reconstructed by adding all the layers, including the colors. Following Tumblin and Turk [62], we also apply a gain, $g_{\mathrm{color}}$, to the color channels. Finally, the resulting image is applied an overall gain $g_{\mathrm{out}}$ and then gamma-corrected [51, Section 2.9] with an exponent of $1/\gamma$. The parameters for this pipeline are organized in Table˜A.6. The implementation uses OpenCV [5], in particular, the guided filter implementation in the extended image processing (ximgproc) submodule.
@@
to the unit hypercube ${[0,1]}^{13}$. In particular, the mapped values on the unit interval $[0,1]$ linearly interpolate the domain of each parameter shown in Table˜A.6. For the radius parameters $r_{1},r_{2},r_{3}$, which are categorical, we naively quantize the domain by rounding the output of the interpolation to the nearest integer.
@@ A.8.2 Peptide design
Table˜A.7 specifies the $T=11$ target bacteria used for the peptide design task from Section˜4. The first seven bacteria are Gram negative bacteria (Objective IDs B1-B7) and the last four (Objective IDs B8-B11) are Gram positive. Table˜A.8 gives the $10$ template amino acid sequences used for the “template constrained" variation of the peptide design task from Section˜4.
@@ A.9 Proof that finding $S^{*}_{D_{s}}$ is NP-hard
In this section, we prove Lemma˜3.1, that finding $S^{*}_{D_{s}}$ Is NP-Hard.
@@
We prove Lemma˜3.1 this by reduction from the well-known Maximum Coverage Problem (MCP), which is NP-Hard.
@@
We do not claim the above proof of Lemma˜3.1 as a novel contribution of this work, as it follows straightforwardly from the fact that the well-known Maximum Coverage Problem is NP-hard. Note that proving Lemma˜3.1 would also follow straightforwardly from the work of Ding et al. [13] who demonstrate that the coverage optimization problem generalizes k-means clustering. Rather than providing additional novel contribution, the above proof of Lemma˜3.1 serves to justify our use of a greedy approximation algorithm (Algorithm˜1) to approximate $S^{*}_{D_{s}}$ on each iteration of MOCOBO.
A.10 Approximation proof for Algorithm˜1
In this section, we prove Theorem˜3.2, that Algorithm˜1 is a $(1-\frac{1}{e})$-Approximation.
@@ A.10 Approximation proof for Algorithm˜1
It follows that from Definition˜A.5 that:
@@
The coverage score $c(S)$ as defined in Definition˜A.3 is monotone, i.e., for any $S\subseteq S^{\prime}\subseteq D_{s}$, we have:
@@
The coverage score $c(S)$ as defined in Definition˜A.3 is submodular, i.e., for any $S\subseteq S^{\prime}\subseteq D_{s}$ and any $\mathbf{x}^{*}\in D_{s}\setminus S^{\prime}$, we have:
@@
Proof of Theorem˜3.2 From Lemma˜A.6 and Lemma˜A.6, we know that $c(S)$ is monotone submodular.
Since $c(S)$ is monotone submodular, and Algorithm˜1 approximates $A^{*}_{D_{s}}$ using greedy submodular optimization, it follows from Lemma˜A.8 that Algorithm˜1 achieves:
@@ 3. Answer: [Yes]
  4. Justification: All stated claims are backed-up with results in Section˜4 and the stated focus/scope of the paper accurately reflects what is discussed throughout the rest of the paper.
@@ 8. Answer: [Yes]
  9. Justification: See Section˜A.4 for discussion of limitations.
@@ 13. Answer: [Yes]
  14. Justification: All theoretical claims can be found in Section˜3.1.2. All sets of assumptions and full proofs of each claim are provided in Section˜3.1.2, Section˜A.9, and Section˜A.10. As far as we are aware, each proof is complete and correct.
@@ 18. Answer: [Yes]
  19. Justification: We provide detailed explanation of how our method works in Section˜3 and all additional required details to reproduce results in Section˜4 and Section˜A.7. Additionally, we provide a link to a public GitHub repository containing the source code used to run our method and produce results provided in this paper. This GitHub repository contains organized code that will allow any reader to run MOCOBO on all tasks in this paper. Additionally, the README in the repository provides detailed instructions to make setting up the proper environment and running the code easy for users.
@@ 23. Answer: [Yes]
  24. Justification: We provide a link to a public GitHub repository containing the source code used to run our method and produce results provided in this paper. This GitHub repository contains organized code that will allow any reader to run MOCOBO on all tasks in this paper. Additionally, the README in the repository provides detailed instructions to make setting up the proper environment and running the code easy for users.
@@ 28. Answer: [Yes]
  29. Justification: All chosen hyper-parameters and implementation details are stated in section 4 and Section˜A.7.
@@ 38. Answer: [Yes]
  39. Justification: All compute details are provided in Section˜A.6.
@@ 48. Answer: [Yes]
  49. Justification: See discussion of broader societal impacts in Section˜A.5.
@@ 58. Answer: [Yes]
  59. Justification: Creators of all assets used to produce all results in this paper are cited in Section˜4. All assets used are open source software or models.